# Robust Hyperbolic Learning with Curvature-Aware Optimization

### v3

@@ Abstract
Hyperbolic deep learning has become a growing research direction in computer vision due to the unique properties afforded by the alternate embedding space. The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings. However, current hyperbolic learning approaches are still prone to overfitting, computationally expensive, and prone to instability, especially when attempting to learn the manifold curvature to adapt to tasks and different datasets. To address these issues, our paper presents a derivation for Riemannian AdamW that helps increase hyperbolic generalization ability. For improved stability, we introduce a novel fine-tunable hyperbolic scaling approach to constrain hyperbolic embeddings and reduce approximation errors. Using this along with our curvature-aware learning schema for Lorentzian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning. Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks achieving state-of-the-art results in two domains and drastically reducing runtime.
Machine Learning, ICML
@@ 1 Introduction
Recently, hyperbolic manifolds have gained attention in deep learning for their ability to model hierarchical and tree-like data structures efficiently. Unlike Euclidean space, hyperbolic space has negative curvature, allowing it to represent exponentially growing distances. This makes these manifolds ideal for tasks like natural language processing, graph representation, and recommendation systems (Peng et al., 2022). By embedding data into hyperbolic space, models can capture complex relationships more effectively, often with fewer parameters (Ganea et al., 2018), leading to improved performance and interpretability. Hyperbolic geometry has also found applications in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation(Bdeir et al., 2024; van Spengler et al., 2023; Khrulkov et al., 2020; Atigh et al., 2022), few-shot learning(Liu et al., 2020), and feature representation (Yang et al., 2023).
@@
Two main derivations of hyperbolic space have emerged as learning frameworks: the hyperboloid or the Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$). Typically, the hyperboloid, or Lorentz space, offers better operational stability, as demonstrated by Mishne et al. (2022) but lacks clear definitions for basic vector operations such as addition and subtraction. To bridge this gap, recent research has focused on defining Lorentzian versions of common deep learning operations, such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024).
@@
However, the use of these hyperbolic operations comes with challenges. Computations in hyperbolic space are more complex and expensive due to the increased mathematical complexity and the lack of optimized CUDA implementations, which can slow down training and drastically increase memory requirements. Additionally, optimization on hyperbolic manifolds can be unstable, especially in low floating-point precision environments. This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius for manifolds Mishne et al. (2022); Guo et al. (2022). Otherwise, performance either suffers or training becomes impossible due to the precision errors causing mathematically undefined operations. The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the representational radius and requires the projection of vectors between these different curvatured manifolds. Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting which heavily impacts their generalization ability.
@@
This work addresses key challenges in the Lorentz model. To combat overfitting, we introduce an AdamW-based optimizer that enhances regularization and generalization in hyperbolic learning. For learning instability, we propose an optimization framework that stabilizes curvature parameter learning. We also present a smoother scaling function to replace weight clipping, ensuring hyperbolic vectors remain within the representational radius. To tackle computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning. Improved stability not only boosts model performance but also enables lower-precision training, enhancing efficiency. Our contributions are four-fold:
@@
    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters.
@@
    We present HCNN+, HyperMAtt, and LHEIR, an application of our proposed optimization scheme for the domains of classical image classification, EEG classification, and hierarchical metric learning respectively.
@@
    We empirically show the effectiveness of our proposed methods in three domains, hierarchical metric learning, Image, and EEG classification, where we achieve state-of-the-art performance in two domains with a significant computational speed-up.
Our code is publicly available at [redacted for double-blind submission].
@@ 2 Related Work
With the success of employing hyperbolic manifolds in NLP models (Zhu et al., 2021; Dhingra et al., 2018; Tifrea et al., 2018) hyperbolic embeddings have extended to the computer vision domain. Initially, many of the works relied on a hybrid architecture, utilizing Euclidean encoders and hyperbolic decoders (Mettes et al., 2023). This was mainly due to the high computational cost of hyperbolic operations in the encoder, as well as the lack of well-defined alternatives for Euclidean operations. However, this trend has begun to shift towards the utilization of fully hyperbolic encoders as can be seen in the hyperbolic Resnets by Bdeir et al. (2024) and van Spengler et al. (2023). Both works offer hyperbolic definitions for 2D convolutional layer, batch normalization layer, and an MLR for the final classification head. Bdeir et al. (2024) even attempt to hybridize the encoder by employing the Lorentz manifold in blocks that exhibit higher output hyperbolicity. While this has led to notable performance improvements, both models suffer from upscaling issues. Attempting to replicate these approaches for larger datasets or bigger architectures becomes much less feasible in terms of time and memory requirements. Instead, our approach places higher focus on efficient components to leverage the beneficial hyperbolic properties of the model while minimizing the memory and computational footprint.
@@
Previous work in hyperbolic spaces has explored various approaches to curvature learning. In their studies, Gu et al. (2018) and Giovanni et al. (2022) achieve this by using a radial parametrization that implicitly models variable curvature embeddings under an explicitly defined, fixed 1-curve manifold. This method enables them to simulate K-curve hyperbolic and spherical operations under constant curvature for the mixed-curve manifold specifically, a combination of the Euclidean, spherical, and Poincaré manifold. Other approaches, such as the one by Kochurov et al. (2020), simply set the curvature to a learnable parameter but do not account for the manifold changes in the Riemannian optimizers. This leads to hyperbolic vectors being updated with mismatched curvatures and others being inaccurately reprojected, resulting in instability and accuracy degradation. Additionally, some methods, like the one by Kim et al. (2023), store all manifold parameters as Euclidean vectors and project them before use. While this approach partially mitigates the issue of mismatched curvature operations, it remains less accurate and more computationally expensive. In comparison, our proposed optimization schema maintains the parameters on the manifold and optimizes them directly by performing the necessary operations to transition between the variable curvature spaces.
@@ 3.1 The Lorentz Space
The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$. There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold. The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ which is an important characteristic to note for the stability issues presented later on. The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$.
We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, as the space dimension $\bm{x_{s}}$. Below are the basic operations presented by the manifold with more complex operations provided in Appendix A.
@@
Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by
$d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$ (1)
We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
@@
Since the Lorentz space is a Riemannian manifold, it is locally Euclidean. This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$. The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
@@ 3.2 Riemannian Optimization Schema
In the GeoOpt hyperbolic learning library, Kochurov et al. (2020) introduces the curvature $c$ of hyperbolic space as a learnable parameter. However, no subsequent work has fully taken advantage of this feature despite the authors presenting ready implementations of hyperbolic operations that account for this curvature. Our empirical testing also reveals that naively learning this parameter often leads to instability and performance decline. We attribute these issues to the simplistic curvature updates in Riemannian optimizers, which neglect to adjust hyperbolic operations, weights, and gradients after updating the curvature.
In the current Riemannian optimization approach, Euclidean gradients and momentums are projected onto the manifold to update hyperbolic parameters. However, since optimization depends on curvature, adjusting it during learning makes previous projections, gradients, and parameters inconsistent with the new geometry. Essentially, the current schema applies updates using the new curvature on vectors defined by the old curvature, leading to mismatches.
To illustrate this, LABEL:fig:curve_inconsistencies shows two manifolds with different curvatures. It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second. Thus, using the new curvature operations on points from the old model is not mathematically sound practice. This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to undefined values during the parameter updates.
One example of this is the distance function Equation 1 which requires the bound ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ to hold. If the new curvature parametrization at time t $K_{t}>K_{t-1}$ then the input into the inverse hyperbolic cosine has a lower bound less than 1, leaving room for the operation to be mathematically undefined.
While GeoOpt does not address this particular optimization challenge, it does implement an ”N-stabilize step”. This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold. However, this approach is done after the update step and does not prevent invalid parameter values during optimization.
@@
To address this, we propose a method that tracks both the old and new curvature parameter values $\bm{\theta}_{K}$ throughout updates. Specifically, we propose a simple approach where all updates are first applied under the old curvature, then the new curvature is derived and finally, the hyperbolic vectors are mapped from the old manifold to the new one.
The previously proposed mapping technique in the N-stabilize step alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability. Instead, we recommend projecting points onto the tangent space at the origin of the old curvature $\overline{\bm{0}}_{t-1}$ using the logarithmic map and then projecting back after the curvature update. Gradients and their momentum can be similarly projected or parallel transported Appendix A) to the tangent space. The entire optimization schema can then be seen in Algorithm 1.
Here, we choose to apply parallel transport, or any similar approximation of it, because it preserves the gradient vectors’ directions which is important for the optimization procedure. We also choose to use the tangent space at the origin as the mapping intermediate since it remains relatively stable when the curvature is updated, as shown in LABEL:fig:curve_inconsistencies. Finally, we prove that the vector norms and their hyperbolic distances to the origin are preserved since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$.
@@
It is important to emphasize that our proposed optimization scheme is compatible with existing curvature learning and meta-learning methods. Rather than being an alternative, it serves as a guideline for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.

### v4

@@ Abstract
Hyperbolic deep learning has become a growing research direction in computer vision due to the unique properties afforded by the alternate embedding space. The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings. However, current hyperbolic learning approaches are still prone to overfitting, computationally expensive, and prone to instability, especially when attempting to learn the manifold curvature to adapt to tasks and different datasets. To address these issues, our paper presents a derivation for Riemannian AdamW that helps increase hyperbolic generalization ability. For improved stability, we introduce a novel fine-tunable hyperbolic scaling approach to constrain hyperbolic embeddings and reduce approximation errors. Using this along with our curvature-aware learning schema for Riemannian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning. Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks while greatly reducing runtime. Our code is publicly available at https://github.com/inboxedshoe/Robust-Hyperbolic-Learning
@@ 1 Introduction
Recently, hyperbolic manifolds have gained attention in deep learning for their ability to model hierarchical and tree-like data structures efficiently. Unlike Euclidean space, hyperbolic space has negative curvature, allowing it to represent exponentially growing distances. This makes these manifolds ideal for tasks like natural language processing, graph representation, and metric learning [33]. By embedding data into hyperbolic space, models can capture complex relationships more effectively, and often with fewer parameters [11]. Hyperbolic geometry has also been applied in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation [3, 40, 18, 2], few-shot learning [24], and feature representation [44].
@@
These works rely on two main derivations of the hyperbolic space, the hyperboloid or Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$). Typically, the hyperboloid offers better operational stability, as demonstrated by Mishne et al. [29] but lacks clear definitions for basic vector operations such as addition and subtraction. To bridge this gap, recent research has focused on defining Lorentzian variant of common deep learning operations, such as the feed-forward layer [6, 8, 11], the convolutional layer [6, 34, 8], and multinomial linear regression (MLR) [3].
@@
However, the use of these hyperbolic operations comes with challenges. Computations in hyperbolic space are more complex and expensive, and the lack of optimized CUDA implementations drastically slows down training and increases memory requirements. Additionally, optimizing hyperbolic parameters such as class prototypes or batchnorm means can be unstable, especially in low precision floating-point environments. This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius of the manifolds [29, 16]. The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the embedding space. Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting, which heavily impacts their generalization ability.
@@
This work addresses key challenges in the Lorentz model. To combat overfitting, we introduce an AdamW-based optimizer that enhances regularization and generalization in hyperbolic learning. For learning instability, we propose an optimization framework that stabilizes the learning of curvature parameters. We also present a smoother scaling function to replace weight clipping, ensuring hyperbolic vectors remain within the representational radius. To address computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning. Improved stability not only boosts model performance but also enables lower-precision training, further enhancing efficiency. Our contributions are then four-fold:
@@
    We propose an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters and a formulation for the Riemannian AdamW
@@
    We present LHEIR, HyperMAtt, and HCNN+ as applications of our proposed optimization scheme for the domains of hierarchical metric learning, EEG classification, and computer vision for image classification and generation, respectively.
@@
    We empirically show the effectiveness of our proposed methods in five domains, hierarchical metric learning, EEG classification, graph embedding, image classification, and image generation to show the effectiveness of our optimizer in different problem settings. We improve performance in all domains with a significant computational speed-up.
@@ 2 Related Work
Initially, many hyperbolic deep learning methods relied on a hybrid model architecture that utilizes Euclidean encoders and hyperbolic decoders [28]. Euclidean encoders avoid the high computational complexity of hyperbolic operations, as well as the lack of well-defined hyperbolic alternatives for Euclidean components. However, this trend has begun to shift towards fully hyperbolic models. Chen et al. [6] propose hyperbolic components for a fully connected linear layer, a graph convolution layer, and an attention layer with the square Lorentzian distance as a similarity metric. For the classification head, they learn class prototypes directly on the hyperbolic manifold and use the same distance metric as a loss estimator, following similar work in the literature [2, 19, 28]. This work was further extended to computer vision by Bdeir et al. [3] and van Spengler et al. [40] with both developing fully hyperbolic ResNets, using Riemannian batch normalization layers that rely on a learnable mean also embedded and optimized in hyperbolic space. Bdeir et al. [3] attempt to alleviate these issues by including a hybrid encoder that only applies the hyperbolic components in blocks that exhibit higher embedding hyperbolicity. Although this has led to notable performance improvements, both models suffer from upscaling issues. Attempting to apply these approaches to larger datasets or larger architectures becomes less feasible in terms of time and memory requirements. Instead, our approach places a greater focus on efficient components to leverage the beneficial hyperbolic properties of the model while minimizing the memory and computational footprint.
@@
Previous work in hyperbolic spaces has explored various approaches to curvature learning. In their studies, Gu et al. [14] and Giovanni et al. [13] achieve this by using a parametrization that implicitly models variable-curvature embeddings under an explicitly defined 1-curve manifold. This method enables them to simulate K-curve hyperbolic and spherical operations under constant curvature. They apply this method on mixed-curve manifold embeddings where every portion of the embedding belongs to either the Euclidean, spherical, or Poincaré manifold. Other approaches, such as the one by Kochurov et al. [20], simply set the curvature to a learnable parameter but do not account for the manifold changes while performing the optimization steps with the Riemannian optimizers. This leads to mathematical inconsistencies when updating the hyperbolic parameters, resulting in instability and accuracy degradation. Additionally, some methods, like the one by Kim et al. [19], store all manifold parameters as Euclidean vectors and project them before use. While this approach partially mitigates the issue of mismatched curvature operations, it requires repeated backpropagation through hyperbolic mappings, making it computationally expensive and more susceptible to projection errors. Other methods [3, 40, 6] use a combination of Euclidean parametrization, and direct hyperbolic parameter learning depending on the component and required precision.
@@ 3.1 The Lorentz Space
The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$. There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold. The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~|~\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=-K,~x_{t}>0\}$} implying a negative curvature of $\frac{-1}{K}$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that $\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}\leq-K$ which is an important characteristic to note for the stability issues presented later on. The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}^{K}=[\sqrt{K},0,\cdots,0]^{T}$. We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector as the space dimension $\bm{x_{s}}$. Below are the basic operations presented by the manifold with more complex operations provided in Appendix˜A in the supplementary material.
@@
[…]
@@
[…]
@@ 3.2 Riemannian Optimization Schema
Many existing hyperbolic models adopt hybrid learning approaches that combine Euclidean parameterization with direct hyperbolic optimization. Our work focuses on the latter. This motivates our reliance on GeoOpt [20], a widely adopted library for Riemannian optimization based on the definitions by Bécigneul and Ganea [4], whose methodology is also reflected in related work. Here, the curvature $K$ of the hyperbolic space is set as a learnable parameter. However, empirically, we find that naively optimizing this curvature often leads to instability and performance degradation. We argue that this stems from a fundamental oversight: Riemannian optimizers update curvature without adjusting dependent hyperbolic operations, weights, or gradients, creating geometric inconsistencies.
Given a hybrid model with both Euclidean and hyperbolic parameters, we define the set of learnable parameters as $\bm{\theta}=[\bm{\theta}_{\mathbb{E}},\bm{\theta}^{K}_{\mathbb{L}},\bm{\theta}_{K}]$, where $\bm{\theta}_{\mathbb{E}}$ are the parameters optimized in Euclidean space, $\bm{\theta}^{K}_{\mathbb{L}}$ are the parameters constrained to a Riemannian manifold $\mathcal{M}$ with curvature $K$ and origin $\overline{\bm{0}}^{K}$, and $\bm{\theta}_{K}$ are learnable curvature parameters for the manifold. When optimizing the curvature, we refer to $\bm{\theta}^{K_{t}}_{\mathbb{L}}$ as the hyperbolic parameters defined on the manifold with the value of the curvature parameter at timestep $t$. It is easy to see here that $\bm{\theta}^{K}_{\mathbb{L}}$ is dependent on $\bm{\theta}_{K}$ since the hyperbolic parameters are defined by their values. Current Riemannian optimization first calculates the Euclidean gradient $\mathcal{G}$ through normal backpropagation. If the parameter we are currently updating is Euclidean ($\bm{\theta}_{\mathbb{E}}$), we perform a typical Euclidean update step. If the parameter being updated is hyperbolic ($\bm{\theta}^{K}_{\mathbb{L}}$), the optimizer projects the gradient onto the tangent space of the parameter $\mathcal{T}_{\bm{\theta}_{\mathbb{L}}}$ in the process egrad2rgrad described in Appendix A. Momentum vectors typically used with optimizers like Adam and SGD are also initialized and updated on the tangent space. Finally, the update step is performed using some form of retraction or exponential map.
Currently, Riemannian optimizers treat $\bm{\theta}_{K}$ as a Euclidean parameter. However, since the curvature defines the geometry of the manifold, changing it during training renders prior projections, gradient momentums, and $\bm{\theta}^{K}_{\mathbb{L}}$ parameters misaligned with the new curvature. GeoOpt attempts to mitigate this instability through an "N-stabilize step," which periodically recomputes the time components of hyperbolic parameters to ensure adherence to the manifold. However, this occurs after updates, failing to prevent invalid intermediate states during optimization.
As a concrete example, let $H$ and $H^{\prime}$ be two Lorentz manifolds with curvatures $K$ and $K^{\prime}$. Let $\mathbf{x}$ be a learnable model parameter that lies on $H$, meaning it satisfies $\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=-K$. The hyperbolic distance from this point to itself, $d_{\mathcal{L}}(\mathbf{x},\mathbf{x})$, should always be 0. However, if the optimizer first updates the curvature to $K^{\prime}$ but does not yet update the parameter $\mathbf{x}$, any subsequent operation will use the new curvature $K^{\prime}$ with the old parameter coordinates. The distance calculation becomes:
$d_{\mathcal{L}}(\mathbf{x},\mathbf{x})=\sqrt{K^{\prime}}\cdot\operatorname{acosh}\left(\frac{-\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}}{K^{\prime}}\right)=\sqrt{K^{\prime}}\cdot\operatorname{acosh}\left(\frac{K}{K^{\prime}}\right)$
If $K<K^{\prime}$, the term $\frac{K}{K^{\prime}}$ is less than 1. The $\operatorname{acosh}$ function is undefined for inputs less than 1, leading to NaN values and a model crash. If $K>K^{\prime}$, the calculation does not crash, but the distance is no longer 0. This incorrect distance creates invalid gradients, which destabilizes training. Previous similar works relied on rigorous training regimes with separate optimizers for the curvature, very small learning rates, and "burn-in" epochs [36, 7] which we attempt to avoid. Other solutions such as clipping may prevent undefined operations but cannot easily account for the inconsistencies in the second scenario.
@@
To address this, we propose a method that groups parameters based on their respective spaces and staggers their updates. Specifically, at timestep $t$ we first isolate $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ and update these parameters first. We then update $\bm{\theta}_{\mathbb{E}}$ followed by $\bm{\theta}_{K}$. At this stage, we have updated all the model parameters, but $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ are still defined on the old curvature, so we must use a mapping function to update $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}\rightarrow\bm{\theta}^{K_{t}}_{\mathbb{L}}$.
The N-stabilize step used by GeoOpt can be seen as a pseudo-map, but it alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability. As such, we identify two alternate mapping techniques commonly used in the literature and compare their properties. We start with the scaling function used by Tabaghi et al. [38], Skopek et al. [36]. In their works the authors simply multiply $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ by the scaling factor $l=\sqrt{\frac{K_{t}}{K_{t-1}}}$; this would automatically map all the parameters from $\mathbb{L}^{n}_{K_{t-1}}$ to $\mathbb{L}^{n}_{K_{t}}$ and scale all distances between points linearly by that factor. As such, it preserves the relative distances between the hyperbolic parameters.
However, the method itself presents issues during optimization and input scaling. Hyperbolic models suffer from instability and traditionally require higher precision computations to prevent invalid values. Mishne et al. [29] show that the mathematical instability is proportional to the distance between $\overline{\bm{0}}$ and the hyperbolic embeddings. They are also able to derive the maximum distance allowed before we begin to get undefined operations. By using the scaling factor $l$ we also scale $d_{\mathbb{L}}(\bm{x},\overline{\bm{0}})$, which could then push it outside the stable range. We theorize this is why works using this method need to adhere to strict training regimes, including burn-in periods without curvature learning, a separate optimizer for the curvature, and very low learning rates. Clipping these parameters is one solution, but larger $l$ values could make clipping too harsh, and it removes the nice property of preserved relative distances.
Additionally, there is the issue of dealing with input projections, almost all hyperbolic learning approaches assume that the Euclidean or input data exists on the tangent plane of the origin and project it onto the manifold using the exponential map. This projection preserves the norm of the Euclidean vectors as $d_{\mathbb{L}}(\bm{x},\overline{\bm{0}})$. When the input is then projected onto the scaled manifold, the relative norms of the input and the hyperbolic class prototypes, for example, are now completely different, breaking their relationship. One would then have to find a way to scale the inputs while adhering to the maximum stability distance without clipping. This inconsistency is the same for mixed optimization settings where both Euclidean trivialized parametrization and direct hyperbolic parameter learning are used.
An alternative mapping method is presented in Fu et al. [10], Guo et al. [15] and is based on projecting the parameters onto the tangent space at the origin $\overline{\bm{0}}_{t-1}$ of the old curvature using the logarithmic map and then projecting back after the curvature update. We show in Appendix D that this method preserves the distances and angles between the hyperbolic parameters and the origin. These properties are important as they are considered proxies for hierarchy level, and embedding similarity. We also show why the tangent space at the origin is the most mathematically suitable space for this mapping. This method also mitigates the instability caused by scaling, which removes the need for a more complicated training regime and maintains symmetric parameter handling by operating analogously to trivialized parameter learning.
@@
Given the above, our work relies on the tangent-based mapping method. We extend it to the optimization process by additionally defining the mapping of parameter gradients and momentums from $\mathcal{T}_{\bm{\theta}^{K_{t-1}}_{\mathbb{L}}}$ to $\mathcal{T}_{\bm{\theta}^{K_{t}}_{\mathbb{L}}}$. The entire mapping schema is shown in Algorithm˜1. We additionally perform empirical experiments to comparing the scaling mapping and the tangent mapping in Appendix E. It is important to emphasize that our proposed optimization scheme is compatible with existing curvature learning and meta-learning methods. Rather than being an alternative, it serves as an intermediate step for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.